# Large Language Models (LLMs) for Wireless Networks: An Overview from the Prompt Engineering Perspective

### v1

@@ III Prompt Engineering for Wireless Networks
This section presents several prompting techniques and network applications, i.e., in-context learning, chain-of-thought, prompt-based planning, and self-refinement methods.
@@
Meanwhile, wireless networks are complicated systems, and one task may include several logic and reasoning steps. For instance, many network optimization problems include several elements, e.g., base stations, transmission power, bandwidth, and network users. Using simple input-output pairs as demonstrations cannot provide the logical reasons behind these examples, and LLMs may have difficulty in learning. To this end, chain-of-thought is proposed as an advanced strategy to enhance LLM’s performance in complex reasoning tasks. Specifically, chain-of-thought will provide explanations for each logic step in the prompt demonstrations: “Let’s think step by step. Given the current base station transmission power $\{BS\_power\}$, using the Shannon capacity equation, the transmission data rate is $\{user\_rate\}$. Then we compare the achieved rate with the target data rate, and $\{user\_rate\}$ is lower. Therefore, we may need to increase the transmission power.” This example includes two steps, “using Shannon capacity equation” and “compare the achieved rate with the target data rate”, then the final decision is to “increase the transmission power”. With these two-step explanations, LLMs can easily capture the relationship between transmission power and target data rate, and generate better replies.
In-context learning and chain-of-thought are mainly designed for single tasks, which may have difficulty tackling more complicated tasks, e.g., collecting information from multiple sources and producing multi-step reasoning to generate a comprehensive answer. Many complex network tasks need to be broken down into more manageable sub-tasks. For example, the network configuration usually involves multiple inter-related network elements, and network project development may consist of a series of coding and testing tasks. In this case, prompt-based planning aims to generate a group of sub-tasks, which will be executed one by one. This planning capability is crucial for handling many large-scale network tasks.
In addition, LLMs may generate incorrect answers at the initial attempts, and external feedback is essential to improve its replies. Self-refinement allows LLMs to self-improve their previous outputs by generating feedback and refinement prompts. In particular, it indicates that LLMs can constantly improve their previous outputs automatically without human intervention. Self-refinement can significantly save human effort in improving LLM’s initial replies, especially considering the complexity of network tasks.
@@ IV Iterative Prompting for Wireless Network Optimization
Optimization techniques are of great importance for wireless network management, which has been extensively studied for various network tasks, e.g., power control, resource allocation, user association, task offloading, etc. Considering its crucial importance, this section proposes a novel iterative prompting scheme for network optimization, in which we prompt LLMs iteratively to improve the performance on optimization tasks.
@@ IV-A Natural Language-based Task Description
Firstly, as shown in Fig. 2, the target task is described by formatted natural language, including task goals, definitions, and extra rules, providing fundamental task information to the LLM. For instance, the task goal describes the problem as a “decision-making task”, and it introduces the set of control variables as “candidate decisions”. Then, it highlights the importance of “environment states”, indicating that the LLM has to consider specific environment variables for decision-making. After that, it introduces the current target task by “here is a new case to solve, and the current environment states is…”. Moreover, the final rules specify that “select from {Decision_set} based on above examples”.
With these formatted task descriptions, the LLM can capture the key elements of the decision-making task, focusing on hidden patterns between environment stats and corresponding decision variables. Such a design can also lower the understanding difficulties caused by professional network knowledge, since general domain LLMs such as GPTs and Llama are not specifically pre-trained for network applications.
Figure 3: Self-refined prompting for network traffic prediction. Here, we present an example of hourly traffic prediction for 2013-11-08 using hourly historical traffic from 2013-11-07. Due to space constraints, we show traffic only for certain hours, while traffic is available each hour during interactions with an LLM. The specific feedback related to initial predictions is incorporated into the feedback prompt. The LLM refines its predictions by following the actionable steps outlined in the refinement prompt. Due to space constraints, we only show feedback generation and prediction refinement at one iteration.
@@ IV-B Iterative Prompting Scheme
As illustrated in Fig. 2, the above task description will become input prompts for the LLM at network edge cloud, and the LLM selects a specific Decision_n from the set of candidate decisions {Decision_set}. Then, the network management decisions will be sent to the wireless network environment for implementation, e.g., resource allocation, user association and transmission power level decisions. The network operation results be collected as a new example, including environment states, selected decisions, and rewards that are related to optimization objectives and network metrics.
After that, the accumulated examples consist of an experience pool, recording all previous explorations and examples. They will be further evaluated to select examples that are related to the target task. For instance, a straightforward strategy is to select previous examples with similar environment states as current states, e.g., similar user numbers and channel conditions. Then the previous decisions and rewards become useful references for LLM decision-making. However, note that many wireless environment states are continuous variables, e.g., channel state information or user-base station distances, and therefore it is unlikely to find examples with exactly the same states. Therefore, the evaluation and selection of examples are crucial in the proposed scheme.
Finally, the selected examples, i.e., recommended and inadvisable demonstrations, are integrated into the task description, providing useful references for LLM decision-making. Fig. 2 has presented steps 1 to 6 of the defined prompting scheme, and it can iteratively explore new decisions and accumulate new experiences. Therefore, the LLM can constantly learn from previous explorations, and then improve its decisions accordingly. Such an iterative prompting method can rapidly adapt to dynamic wireless environments. Meanwhile, the proposed scheme in Fig. 2 has several advantages over conventional machine learning approaches: 1) No model parameter updating and fine-tuning is required in this scheme, which has much lower complexity than existing algorithms such as deep reinforcement learning (DRL); 2) It utilizes natural language for network management, making network control and optimization more accessible with minimum requirements on mathematical knowledge and equations; 3) LLMs can provide reasonable explanations for their decisions, and help humans understand complicated network systems, while non-previous algorithms have such ground-breaking capabilities.
@@ V-B Feedback Generation and Prediction Refinement
To summarize, the proposed prompting-based prediction method can constantly improve its prediction performance with self-refinement. Compared with conventional techniques such as long short-term memory (LSTM) networks, the prompting-based approach avoids the complexity of dedicated model training for specific datasets. It is regarded as a plug-and-predict method, providing instant prediction results without parameter fine-tuning. However, the iterative appending of all historical predictions may increases the length of demonstration prompts, potentially leading to higher computational costs and resource consumption.
@@
(c) Prediction Case Study: average MAE and MSE comparison.
(d) Prediction Case Study: predicted traffic comparison.
@@ VI-A Simulation Settings
This section presents two case studies on network optimization and prediction tasks: 1) Network Optimization: This case study involves a base station power control problem, which is a fundamental optimization task that has been widely investigated in wireless network fields. The considered problem formulations can be easily found in [13], which aims to minimize the base station power consumption and meanwhile maintain an average data rate threshold for users. We consider three adjacent base stations and the associated user numbers dynamically change from 5 to 15 each, the average data rate threshold is 1.5 Mbps/per user, and the channel gain applies 3GPP urban network models. ① Iterative prompting: The simulation includes 2 LLM models: Llama3-7b-instruct as a small-scale model, and Llama3-70b-instruct as a large-scale model. ② DRL baseline: DRL is included as an optimal baseline, which has been widely used to address network optimization problems.
2) Network Prediction: This case study considers a well-known network traffic dataset from the city of Milan [14]. It records the time of user interactions and the base stations managing these interactions. Each base station covers an area of approximately $1km\times 1km$, resulting in a total of 8,923 samples. ① Self-refinement: Our proposed self-refined prompting utilizes GPT-4 as the foundation model for one-day-ahead traffic prediction using historical traffic from the previous day. ② LLM-related baselines: Meanwhile, we include GPT-4 and GPT-3.5 without self-refinement as two LLM-related baselines, better demonstrating the capabilities of self-refinement. ③ Conventional baselines: We also consider autoregressive integrated moving average (ARIMA) and LSTM as two conventional baselines. We deploy LSTM as an optimal baseline, since it has been specifically trained on the target network traffic dataset. By contrast, ① Self-refinemen and ② LLM-related baselines have no previous knowledge on the target task.
@@ VI-B Simulation Results and Analyses
1) Network Optimization Case Study: Fig. 4(a) and 4(b) present the optimization task results. Specifically, one can observe that the average power consumption is constantly improved with the increasing number of episodes in Fig. 4(a). It indicates that the proposed scheme in Section IV enables LLMs to learn from previous examples and experience, and then improve its performance on target tasks iteratively. Fig. 4(a) also shows that LLMs present comparable performance as the DRL algorithm, which is achieved without model training or fine-tuning as in conventional ML algorithms. Meanwhile, Fig. 4(a) compares the service quality of Llama3-7b, Llama3-70b, and DRL, which means the probability of violating the preset average data rate threshold of 1.5 Mbps/per user. It highlights that LLMs can maintain high service quality at the beginning of the training process, while the DRL algorithm has a lower service quality in the initial exploration phase. Fig. 4(a) reveals the potential advantage of LLMs in terms of learning efficiency. Given the dedicated designed task description and examples, LLM can quickly learn from human instructions and previous demonstrations. By contrast, the DRL approach relies on exploration and value updating to understand the task, and the low sampling efficiency is a well-known issue of DRL methods.
2) Network Prediction Case Study: Fig. 4(c) compares the average MAE and MSE metrics of various techniques. It is observed that the self-refined prompting method significantly outperforms ARIMA, GPT-3.5, and GPT-4, demonstrating the capabilities of our proposed self-refined prediction method. The self-refined prompting method exhibits notable improvements over GPT-3.5 and GPT-4, with MAE reductions of 29.72% and 17.09%, which underscores more powerful generalization capability on unseen datasets through iterative feedback generation and prediction refinement processes. GPT-3.5 shows a worse performance due to its outdated architecture and designs, which have been observed in many existing studies. In addition, LSTM shows the best performance as an optimal baseline. However, it is worth noting that LSTM has been specifically trained on the target dataset, while all LLM-based methods are zero-shot predictors. Furthermore, similar results can be observed in Fig. 4(d), which illustrates the one-day ground truth traffic and the predicted traffic of various methods. The self-refined prediction method more closely matches the ground truth traffic, indicating strong scalability of predictions across various base stations. By contrast, GPT-4 shows an obvious mismatch from 9:00 to 15:00.
@@
VII Conclusion
The progress of generative AI and LLMs brings promising opportunities for next-generation wireless networks. This work provides a comprehensive overview of prompt engineering for network applications. It presents in-depth analyses of different prompting techniques such as in-context learning, chain-of-thought, and self-refinement. In addition, this work proposes two novel prompting schemes for network optimization and prediction problems, and the case studies show that the proposed methods can achieve satisfactory performance.

### v2

@@ III Prompt Engineering for Wireless Networks
This section presents several prompting techniques and their network applications, i.e., in-context learning, chain-of-thought, prompt-based planning, and self-refinement methods.
@@
Meanwhile, wireless networks are complicated systems, and one task may include several logic and reasoning steps. For instance, many network optimization problems include several elements, e.g., base stations, transmission power, bandwidth, and network users. Using simple input-output pairs as demonstrations cannot provide the logical reasons behind these examples, and LLMs may have difficulty in learning. To this end, chain-of-thought is proposed as an advanced strategy to enhance LLM’s performance in complex reasoning tasks. Specifically, chain-of-thought will explain each logic step in the prompt demonstrations: “Let’s think step by step. Given the current base station transmission power $\{BS\_power\}$, using the Shannon capacity equation, the transmission data rate is $\{user\_rate\}$. Then, we compare the achieved rate with the target data rate, and $\{user\_rate\}$ is lower. Therefore, we may need to increase the transmission power.” This example includes two steps, “using Shannon capacity equation” and “compare the achieved rate with the target data rate”, and then the final decision is to “increase the transmission power”. With these two-step explanations, LLMs can easily capture the relationship between transmission power and target data rate, and generate better replies using the logical chain.
In-context learning and chain-of-thought are mainly designed for single tasks, which may have difficulty tackling more complicated tasks, e.g., collecting information from multiple sources and producing multi-step reasoning to generate a comprehensive answer. Many complex network tasks need to be broken down into more manageable sub-tasks. For example, the network configuration usually involves multiple interrelated network elements, and network project development may consist of a series of coding and testing tasks. In this case, prompt-based planning aims to generate a group of sub-tasks, which will be executed one by one. This planning capability is crucial for handling many large-scale network tasks.
In addition, LLMs may generate incorrect answers at the initial attempts, and external feedback is essential to improve its replies. As shown in Fig.1, self-refinement allows LLMs to self-improve their previous outputs by generating feedback and refinement prompts. For instance, the LLM’s initial transmission power decision may be incorrect, and we can use the same LLM to evaluate and provide feedback, and then feed the feedback and refinement to the LLM, improving the initial power control decisions. In particular, it indicates that LLMs can automatically improve their previous outputs without human intervention. Self-refinement can significantly save human effort in improving LLM’s initial replies, especially considering the complexity of network tasks.
@@ IV Iterative Prompting for Wireless Network Optimization
Optimization techniques are of great importance for wireless network management, and this section proposes a novel iterative prompting scheme, in which we prompt LLMs iteratively to improve the performance on target network tasks.
@@ IV-A Natural Language-based Task Description
Firstly, as shown in Fig. 2, the target task is described by formatted natural language, including task goals, definitions, and extra rules, providing fundamental task information to the LLM. For instance, the task goal describes the problem as a “decision-making task”, and it introduces the set of control variables as “candidate decisions”. Then, it highlights the importance of “environment states”, indicating that the LLM has to consider specific environment variables for decision-making. After that, it introduces the current target task by “here is a new case to solve, and the current environment state is…”. Moreover, the final rules specify that “select from {Decision_set} based on above examples”.
With these formatted task descriptions, the LLM can capture the key elements of the decision-making task, focusing on hidden patterns between environment states and corresponding decision variables. Such a design can also lower the understanding difficulties caused by professional network knowledge, since general domain LLMs such as GPTs and Llama are not specifically pre-trained for network applications.
Figure 3: Self-refined prompting for network traffic prediction. As an example, by following prediction tasks framed in a question-answer format, the LLM is guided to predict hourly traffic prediction for 2013-11-08, using hourly historical traffic from 2013-11-07. Due to space constraints, we show traffic for selected hours only, while hourly traffic is fully utilized during interactions with the LLM. The specific feedback related to initial predictions is incorporated into the feedback prompt. The LLM further refines its predictions by following the actionable steps outlined in the refinement prompt. The process of feedback generation and prediction refinement is iteratively conducted until the prediction performance converges, whereas the inference is completed without engaging in the feedback generation and prediction refinement.
@@ IV-B Iterative Prompting Scheme
As illustrated in Fig. 2, the above task description will become input prompts for the LLM at network edge cloud, and the LLM selects a specific Decision_n from the set of candidate decisions {Decision_set}. Then, the network management decisions will be sent to the wireless network environment for implementation such as resource allocation, user association and transmission power level decisions. The network operation results be collected as a new example, including environment states, selected decisions, and rewards that are related to optimization objectives and network metrics.
After that, the examples from the network output, i.e., previous explorations, are collected in an experience pool, recording all previous examples. The examples in this pool will be further evaluated to choose specific examples well-suited for the target task. For instance, a straightforward strategy is to select previous examples with similar environment states as current states, e.g., similar user numbers and channel conditions. Then the previous decisions and rewards become useful references for LLM decision-making. However, note that many wireless environment states are continuous variables, e.g., channel state information or user-base station distances, and therefore it is unlikely to find examples with exactly the same states. Therefore, the evaluation and selection of examples are crucial in the proposed scheme.
Finally, the selected examples, i.e., recommended and inadvisable demonstrations, are integrated into the task description, providing useful references for LLM decision-making. Fig. 2 has presented steps 1 to 6 of the proposed iterative prompting scheme, and it can iteratively explore new decisions and accumulate new experiences. Therefore, the LLM can constantly learn from previous explorations, and then improve its decisions accordingly. Such an iterative prompting method can rapidly adapt to dynamic wireless environments.
Fig. 2 applies an exploration-exploitation strategy as in reinforcement learning (RL). However, iterative prompting has several advantages over RL approaches: 1) No model parameter updating and fine-tuning is required, which has much lower complexity than existing RL algorithms such as deep reinforcement learning (DRL). 2) It utilizes human natural language for network management. With minimum requirements on mathematical knowledge and equations, it makes network optimization more accessible than using RL. 3) LLMs can provide reasonable explanations for their decisions, and help humans understand complicated network systems, which is beyond the capabilities of RL techniques. Finally, note that our proposed schemes can also be generalized to many other fundamental network optimization tasks, providing a novel LLM-based solution for network control.
@@ V-B Feedback Generation and Prediction Refinement
In summary, LLMs excel at recognizing and performing new tasks through prompt engineering techniques that utilize contextual information, including task descriptions and demonstration examples. Specifically, the proposed self-refined prompting method designs task-specific demonstration, feedback, and refinement prompts, allowing the LLM to iteratively enhance its predictions by adhering to the instructions in these prompts, without requiring additional model fine-tuning or training. Compared to the dedicated training of long short-term memory (LSTM) networks, the proposed self-refined prompting method is more computationally efficient with the inference abilities, particularly advantageous for on-device learning on resource-constrained devices in real-world wireless networks. Given that iteratively appending all historical predictions can increase the length of demonstration prompts and potentially lead to higher resource consumption, a future direction for the self-refined prompting method will focus on effectively compressing extensive demonstration prompts, thereby further enhancing resource efficiency.
@@
(c) Prediction Case Study: average MAE and MSE comparison. The self-refine method achieves comparable performance as LSTM as a few-shot predictor, while LSTM has been delicately trained on the dataset.
(d) Prediction Case Study: predicted traffic comparison in 24 hours.
@@ VI-A Simulation Settings
This section presents two case studies on network optimization and prediction problems.
1) Network Optimization: This case study involves a base station power control problem a fundamental optimization task in wireless network fields. The considered problem formulations aim to minimize the base station power consumption and meanwhile maintain an average data rate threshold for users [^1] Detailed equations can be found in [13] (equation (1.12) in Section 1.3.5). We consider three adjacent base stations and the associated user numbers dynamically change from 5 to 15 each, the average data rate threshold is 1.5 Mbps/per user, and the channel gain applies 3GPP urban network models. This case study includes: ① Iterative prompting: We consider 2 LLM models: Llama3-7b-instruct as a small-scale model, and Llama3-70b-instruct as a large-scale model. We deploy the proposed iterative prompting technique as introduced in Fig. 2 and Section IV, in which the LLM will explore the environment, accumulate experience, and learn iteratively. ② DRL baseline: DRL is included as a baseline, which has been widely used to address network optimization problems after dedicated model training. We apply a classic deep Q-learning algorithm, in which the neural networks have 3 layers, the experience pool size is 10000, the batch size is 64, and the learning rate is 0.005.
2) Network Prediction: The evaluation dataset is a publicly available network traffic dataset from the city of Milan [14], recording the time of user interactions and the base stations managing these interactions. The original grid of $100\times 100$ base stations is reorganized into $25\times 25=625$ aggregated base stations. Each aggregated base station covers an area of approximately $1km\times 1km$, including 8,923 samples. For the evaluation, we randomly select one base station, partitioning the dataset such that 70% is used for training, 10% for validation, and the remaining 20% for testing. The simulation algorithms include: ① Self-refinement: Our proposed self-refined prompting utilizes GPT-4 as the foundation model for one-day-ahead traffic prediction using historical traffic from the previous day. It follows the proposed self-refined prompting techniques as defined in Fig. 3 and Section V. ② LLM-related baselines: We include GPT-4 and GPT-3.5 without self-refinement as two LLM-related baselines, better demonstrating the capabilities of self-refinement. ③ Conventional baselines: Autoregressive integrated moving average (ARIMA) and LSTM as two conventional baselines. Through extensive experiments, the optimal parameters for the ARIMA baseline model are determined including autoregressive ($p=2$), differences ($d=1$), and moving average ($q=2$) components. We deploy LSTM as an optimal baseline, since it has been specifically trained on the target network traffic dataset. By contrast, Self-refinement and LLM-related baselines have no previous knowledge on the target task.
@@ VI-B Simulation Results and Analyses
1) Network Optimization Case Study: Fig. 4(a) and 4(b) present the optimization task results. Specifically, one can observe that the average power consumption is constantly improved with the increasing number of episodes in Fig. 4(a). It indicates that the proposed scheme in Section IV enables LLMs to learn from previous examples and experience, and then improve its performance on target tasks iteratively. Fig. 4(a) also shows that LLMs present comparable performance as the DRL algorithm, which is achieved without model training or fine-tuning as in conventional ML algorithms. Meanwhile, Fig. 4(b) compares the service quality of Llama3-7b, Llama3-70b, and DRL, which means the probability of violating the preset average data rate threshold of 1.5 Mbps/per user. It highlights that LLMs can maintain high service quality at the beginning of the training process, while the DRL algorithm has a lower service quality in the initial exploration phase. Fig. 4(b) reveals that LLMs have a higher learning efficiency than the DRL algorithm. It can be explained by LLM’s few-shot learning capabilities, in which LLMs can quickly learn from demonstrations and task descriptions to improve decisions. By contrast, the DRL approach relies on exploration and value updating to improve the policy gradually, and the low sampling efficiency prevents the learning efficiency.
2) Network Prediction Case Study: Fig. 4(c) compares the average MAE and MSE metrics of various techniques. It is observed that the self-refined prompting method significantly outperforms ARIMA, GPT-3.5, and GPT-4, demonstrating the capabilities of our proposed self-refined prediction method. The self-refined prompting method exhibits notable improvements over GPT-3.5 and GPT-4, with MAE reductions of 29.72% and 17.09%, which underscores more powerful generalization capability on unseen datasets through iterative feedback generation and prediction refinement processes. GPT-3.5 shows a worse performance due to its outdated architecture and designs, which have been observed in many existing studies. Although LSTM serves as an optimal baseline with superior performance, it is important to note that LSTM is specifically trained on the target dataset, incurring additional computational costs, which presents significant challenges in scenarios involving resource-constrained devices within wireless networks. Given the dynamic evolution of traffic distribution in non-stationary wireless networks, a well-trained LSTM may overfit the historical traffic used during the pre-training, resulting in degraded performance on new, unseen traffic over time. Furthermore, similar results can be observed in Fig. 4(d), which illustrates the one-day ground truth traffic and the predicted traffic of various methods. The self-refined prediction method more closely matches the ground truth traffic, indicating strong scalability of predictions across various base stations. By contrast, GPT-4 shows an obvious mismatch from 9:00 to 15:00.
@@
VII Real-world Challenges and Limitations
This section will further discuss the limitations and real-world challenges of prompt engineering applications.
Firstly, prompt engineering is inherently limited by the context window size of LLMs, which restricts the capabilities to process tasks requiring extensive contextual information beyond the model’s input and memory constraints. Recent advancements such as Llama3 with a context window size of 128k tokens, significantly mitigate this limitation.
Secondly, the effectiveness of prompt engineering relies heavily on the selection of demonstrations shown to the model or the order of demonstrations. Poorly chosen or suboptimal examples can result in biased outputs. Our work considers iterative prompting to select better demonstrations, but professional experiences are still needed in this process.
Moreover, crafting effective prompts for complicated tasks in network environments usually requires trial-and-error experimentation. This iterative process is inefficient, especially when results are inconsistent or difficult to reproduce. Without clear guidelines, engineers must rely on intuition or extensive testing to achieve desired outcomes. This can slow down the development for novel and complex use cases as in 6G.
In addition, in real-world applications, prompt engineering may introduce significant security concerns, as it is potentially vulnerable to attacks such as prompt injection and leakage. Such attacks are particularly concerning in sensitive domains like wireless communications, where malicious actors could manipulate the model or extract encoded sensitive data[15].
Finally, note that prompt engineering is not a standalone solution but one part of pipelines of LLM and generative AI techniques [5]. Prompt engineering is constrained by the inherent limitations of the used LLMs. If the LLM lacks sufficient knowledge or context, prompting is unlikely to produce accurate results. Therefore, when prompt engineering struggles with highly specific tasks, it may require external support such as fine-tuning. This dependency highlights that prompt engineering can be actively combined with other techniques such as pre-training and fine-tuning to make the most of LLM potential.
VIII Conclusion
The progress of generative AI and LLMs brings promising opportunities for next-generation wireless networks. This work provides a comprehensive overview of prompt engineering for network applications. It presents in-depth analyses of different prompting techniques such as in-context learning, chain-of-thought, and self-refinement. In addition, this work proposes two novel prompting schemes for network optimization and prediction problems, and the case studies show that the proposed methods can achieve satisfactory performance. In the future, we will explore more complicated and advanced network tasks such as reconfigurable intelligence surfaces and integrated sensing and communication, uncovering the full potential of LLMs for future wireless networks.